OpenStudy (anonymous):

Use the quadratic formula to solve (x − 4)(x + 5) = 7

OpenStudy (shayaan_mustafa):

Hi again :)

OpenStudy (anonymous):

lol hello:)

OpenStudy (anonymous):

i think we should be able to ask multiple questions on here at a time because i have quite a few xD

OpenStudy (campbell_st):

expand and collect like terms x^2 +x - 27 = 0 a = 1, b = 1 and c = -27 substitute into \[x = \frac{-b \pm \sqrt{b^2 - 4ac}}{2a}\]

OpenStudy (shayaan_mustafa):

First multiply the the brackets like this, \(\Large x^2+5x-4x-20=7\) \(\Large x^2+x-20=7\) Now it is the quadratic equation. The quadratic equation is the one which has 2nd degree.